The Terrence Co. manufactures two products, Baubles and Trinkets. The following are projections for the coming year:

Baubles Trinkets
12,000 units 6,000 units
Sales $ 12,000 $ 12,000
Costs:
Fixed $ 2,900 $ 6,760
Variable 7,200 10,100 3,000 9,760
Income before taxes $ 1,900 $ 2,240


How many Baubles will be sold at the break-even point, assuming that the facilities are jointly used with the sales mix remaining constant?

rev: 10_26_2018_QC_CS-145247

A. 7,250

B. 8,400

C. 11,270

D. 14,490
Answer #1
Baubles Trinkets
Sales units 12000 6000
Sales mx percentage 66.67% 33.33%
Sales 12000 12000
Less: Variable costs 7200 3000
Contribution margin 4800 9000
Contribution margin per unit 0.40 1.50
Weighted average Contribution margin per unit=(0.40*66.67%)+(1.50*33.33%)=$0.7667
Overall break even point = (2900+6760)/0.7667= 12600 units
Baubles units sold at the break-even point = 12600*66.67% = 8400
Option B 8,400 is correct